The probabilities that a surgeon operates on 3 4 5 6 or 7 patents and any one day are 0.15 0.20 0.25 0.20 and 0.20 respectively


1
Expert's answer
2022-03-10T02:46:50-0500

we define the following

mean= ( (3*0.15) +(4*0.20) +(5*0.25)+(6*0.20)+(7*0.20) ) = 5.1


variance= ( (32 * 0.15) + ( 42 * 0.20) + ( 52 * 0.25) + (62 * 0.20) + (72 * 0.20) ) - (5.12)

=27.8 - 26.01 = 1.79


standard deviation = (variance)1/2 = (1.79)1/2 = 1.337908816